### v1

# Optimizing a-DCF for Spoofing-Robust Speaker Verification

Abstract
Automatic speaker verification (ASV) systems are vulnerable to spoofing attacks such as text-to-speech. In this study, we propose a novel spoofing-robust ASV back-end classifier, optimized directly for the recently introduced, architecture-agnostic detection cost function (a-DCF). We combine a-DCF and binary cross-entropy (BCE) losses to optimize the network weights, combined by a novel, straightforward detection threshold optimization technique. Experiments on the ASVspoof2019 database demonstrate considerable improvement over the baseline optimized using BCE only (from minimum a-DCF of $0.1445$ to $0.1254$), representing $13\%$ relative improvement. These initial promising results demonstrate that it is possible to adjust an ASV system to find appropriate balance across the contradicting aims of user convenience and security against adversaries.
Index Terms:
a-DCF, spoofing-robust speaker verification
I Introduction
Automatic speaker verification (ASV) seeks to verify a claimed identity based on speech evidence [1]. Despite being a matured biometric technology, ASV is susceptible to spoofing through replay [2], text-to-speech [3] and other adversaries. To counter these risks, various countermeasures (CMs) or presentation attack detectors (PADs) [4], have been proposed for detecting spoofing. Typically CMs are developed in isolation from ASV, and the two subsystems are combined to enhance robustness of the biometric system against adversaries. The common strategies include tandem (cascade) combination [5] and fusion of detection scores [6] or embeddings [7, 8]. Besides systems constructed from combination of speaker and spoof detection subsystems, other studies consider deep end-to-end networks optimized to predict the decision directly from the speech inputs [9, 10].
Whether realized by combining two subsystems (ASV and CM), or by training a monolithic end-to-end model, a key consideration is how to evaluate and optimize the complete biometric authentication system. Any such system must trade-off between misses (false rejections) and false alarms (false acceptances)—proxies for user (in)convenience and security, respectively. The so-called tandem detection cost function (t-DCF) [5] can be used to evaluate the performance of cascaded combinations of ASV and CM under a Bayes-risk framework. Nonetheless, as t-DCF requires two sets of detection scores (and two detection thresholds) corresponding to the ASV and the CM subsystems, it is inapplicable for performance assessment of other types of architectures. To address these limitations, [11] proposed recently an architecture-agnostic detection cost function (a-DCF) applicable to a wider variety of system architectures. Unlike t-DCF, a-DCF requires a single detection score and a single threshold value only. This new metric is the primary metric in one of the two main tracks in the currently running ASVspoof 5 challenge [^1] https://www.asvspoof.org/ (referred June 28, 2024).
For the first time, we propose a spoofing-robust ASV back-end that is directly optimized for the new a-DCF metric (as shown in Fig. 1). As an evaluation metric based on hard error counting, the original a-DCF metric is non-differentiable and cannot be directly optimized using gradient descent schemes. Luckily, hard error counts can be “softened” (made differentiable). While direct optimization of the t-DCF was addressed in [12] through reinforcement learning, this approach is limited to the tandem architectures. By integrating a differentiable version of the a-DCF into model training, we aim to enhance the performance and robustness of speaker verification systems against spoofing attacks, ultimately contributing to the advancement of secure and reliable biometric authentication technologies. Besides integrating the a-DCF metric into model training, the threshold value is also jointly optimized during model training in the proposed method.
Figure 1: Proposed a-DCF optimization method for spoofing-robust speaker verification. The model $g(x)$ takes embeddings as input and outputs score distributions, from which error rates $\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}$, $\hat{P}_{\mathrm{fa}}^{\mathrm{non}}$, and $\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}$ are calculated. These error rates are then used to compute soft a-DCF, which serve as the loss function for model optimization through backpropagation, ensuring improved performance by minimizing the a-DCF.
II Detection Cost Functions
Conventional detection cost functions [13] are evaluation metrics used to assess the performance of binary classification systems, such as speaker verification and anti-spoofing. They focus on balancing the trade-off between false alarms and misses, assigning specific costs to each type of error. A binary classifier can be defined as a function $g:\mathcal{X}\to\rm I\!R$ that assigns a detection score $g(x)$ for a given input $x\in\mathcal{X}$. This score then can be converted into predicted binary class label $\hat{y}\in\{0,1\}$ by comparing it with a pre-set threshold $\tau$ by $\hat{y}=\mathrm{I}(g(x)>\tau)$. Here, $\mathrm{I}$ is the indicator function which returns $1$ if the condition is true (null hypothesis accepted) and 00 otherwise. In conventional ASV, $x$ consists of a pair of enrollment and test utterances (presented typically as a pair of speaker embeddings), and the null hypothesis is that the same speaker is present in both. In our spoofing-aware ASV models, $x$ additionally contains spoof embedding extracted from the test utterance by a CM model.
II-A DCF and a-DCF
Detection Cost Function (DCF) [13] is used for performance assessment of conventional ASV systems. It is defined as:
$\mathrm{DCF(\tau_{asv})}=C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{asv}})+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{asv}}),$ (1)
where $C_{\mathrm{miss}}^{\mathrm{tar}}$ is the cost of a miss, $C_{\mathrm{fa}}^{\mathrm{non}}$ is the cost of a false alarm, $P_{\mathrm{miss}}^{\mathrm{tar}}$ is the miss rate, $P_{\mathrm{fa}}^{\mathrm{non}}$ is false alarm rate, $\pi_{\mathrm{tar}}$ is the prior probability of the target and $\pi_{\mathrm{non}}=1-\pi_{\mathrm{tar}}$ is the prior probability of nontarget, and $\tau_{\mathrm{asv}}$ is the detection threshold. The DCF in Eq. (1) allows one to set the cost parameters according to the user’s application specification.
The architecture-agnostic detection cost function (a-DCF) [11] is a performance metric introduced for spoofing-aware speaker verification (SASV) task. It extends upon the DCF in Eq. (1) by including a third class (spoofing attack):
$\mathrm{a\!-\!DCF}(\tau_{\mathrm{sasv}})=$ $\;C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ (2)
$+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$
$+C_{\mathrm{fa}}^{\mathrm{spf}}\cdot\pi_{\mathrm{spf}}\cdot P_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}}),$
where $P_{\mathrm{fa}}^{\mathrm{spf}}$ is the false alarm for spoof trials, $\pi_{\mathrm{spf}}$ is the prior probability of spoofing, $C_{\mathrm{fa}}^{\mathrm{spf}}$ is the respective costs for spoof false alarms. $\tau_{\mathrm{sasv}}$ is the detection threshold for spoofing-aware speaker verification (SASV) system.
The miss rate ($P_{\mathrm{miss}}^{\mathrm{tar}}$) and the false alarm rates ($P_{\mathrm{fa}}^{\mathrm{non}}$, $P_{\mathrm{fa}}^{\mathrm{spf}}$) in the a-DCF metric are computed as follows:
$P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{tar}}}\sum_{x\in\mathrm{tar}}\mathrm{I}(g(x)\leq\tau_{\mathrm{sasv}})$ (3)
$P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{non}}}\sum_{x\in\mathrm{non}}\mathrm{I}(g(x)>\tau_{\mathrm{sasv}})$
$P_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\text{sasv}})$ $=\frac{1}{N_{\mathrm{spf}}}\sum_{x\in\mathrm{spf}}\mathrm{I}(g(x)>\tau_{\mathrm{sasv}}),$
where $N_{\mathrm{tar}}$, $N_{\mathrm{non}}$, and $N_{\mathrm{spf}}$ represent the number of target, non-target, and spoof class trials, respectively.
II-B Differentiable a-DCF Metric
The indicator function used to compute false alarm and miss rates in Eq. (3) is non-differentiable. Therefore, the a-DCF metric can not be directly used as loss function to optimize an SASV system. In this study, we propose a differentiable (soft) version of a-DCF suitable for direct minimization using gradient descent. Similar approaches have been successfully applied to optimize DCF for text-dependent speaker verification [14] and t-DCF for the constrained class of tandem systems [15], but not the a-DCF. The a-DCF metric provides a more holistic evaluation by integrating the costs associated with speaker verification. This approach simplifies the benchmarking process for different ASV system architectures, as it eliminates the need for separate metrics for each component. By focusing on a single metric, the a-DCF metric promotes development of integrated systems that are optimized for spoofing-aware speaker verification.
The non-differentiable components in a-DCF metric are the miss rate ($P_{\text{miss}}$) and the two false alarm rates ($P_{\text{fa}}^{\text{non}}$, $P_{\text{fa}}^{\text{spf}}$). A common approach for ‘softening’ hard error counts [14, 15] is to replace the indicator function by a continuous, differentiable function. We adopt the following approximations:
$\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{tar}}}\sum_{x\in\mathrm{tar}}\sigma\left(\tau_{\mathrm{sasv}}-g(x)\right)$ (4)
$\hat{P}_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{non}}}\sum_{x\in\mathrm{non}}\sigma\left(g(x)-\tau_{\mathrm{sasv}}\right)$
$\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{spf}}}\sum_{x\in\mathrm{spf}}\sigma\left(g(x)-\tau_{\mathrm{sasv}}\right),$
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function. Essentially, the error rates are approximated based on their distance from the threshold. The soft a-DCF loss is now defined as:
$\mathcal{L}_{\mathrm{a\!-\!DCF}}^{\mathrm{soft}}(\tau_{\mathrm{sasv}})=$ $\;C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ (5)
$+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot\hat{P}_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$
$+C_{\mathrm{fa}}^{\mathrm{spf}}\cdot\pi_{\mathrm{spf}}\cdot\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}}).$
II-C Optimization of soft a-DCF
Algorithm 1 Optimization Algorithm
1:Input: Train data $\mathcal{D}_{\mathrm{trn}}$, Development data $\mathcal{D}_{\mathrm{dev}}$, Batch Size $B$, Number of Epochs $N$
2:Output: Best threshold $\hat{\tau}_{*}$, Model Parameters $\Theta_{*}$
3:Initialize: $\mathrm{min\_a\!-\!DCF}\leftarrow\infty$, $\tau\leftarrow 0.5$
4:for epoch $=1$ to $N$ do
5:    // optimize DNN weights
6:    for $X\leftarrow get\_minibatch(\mathcal{D}_{\mathrm{trn}},B)$ do
7:         $\mathcal{J}(\Theta_{i},\tau_{i})\leftarrow\left(\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(X;\Theta_{i},\tau_{i})+\mathcal{L}_{\mathrm{BCE}}(X;\Theta_{i})\right)/2$
8:         $\Theta_{i+1}\leftarrow update\_network\_weights(\Theta_{i},\mathcal{J}(\Theta_{i},\tau_{i}))$
9:    end for
10:    $\Theta\leftarrow\Theta_{i+1}$
11:    // find optimal threshold using grid search
12:    $\tau_{i+1}\leftarrow\operatorname*{argmin}\limits_{\tau}\{\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\Theta,\tau)\}$ in $\mathcal{D}_{\mathrm{trn}}$
13:    // keep track of best model using dev data
14:    if $\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\tau_{i+1})<\mathrm{min\_a\!-\!DCF}$ in $\mathcal{D}_{\mathrm{dev}}$ then
15:         $\mathrm{min\_a\!-\!DCF}\leftarrow\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\tau_{i+1})$
16:         $\hat{\tau}\leftarrow\tau_{i+1}$
17:    end if
18:end for
19:return $\hat{\tau}_{*}$, $\Theta_{*}$
The algorithm 1 presents the training loop of a neural network towards its optimization, with both soft a-DCF ($\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(X;\Theta_{i},\tau_{i})$ ) and BCE loss ( $\mathcal{L}_{\mathrm{BCE}}(X;\Theta_{i})$ ) functions. The BCE loss is the binary cross entropy loss defined as:
$\mathcal{L}_{\mathrm{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}[y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})]$ (6)
where $N$ represents the number of training samples. Essentially, the soft a-DCF loss is a loss function of application dependent a-DCF metric and useful for fine-tuning the system towards a desired miss-false alarm balance. The BCE loss, in turn, helps to separate target from the pool of the two negative classes. For these reasons, we use the combination of these two losses. The algorithm initializes a minimum a-DCF to be infinity and a fixed threshold $\tau$. Then, for each epoch, it optimizes the model weights by calculating the combined loss $\mathcal{J}(\Theta_{i},\tau_{i})$ from soft a-DCF and BCE losses and updating the weights of the neural network. This combined loss is defined as:
$J(\Theta_{i},\tau_{i})=\frac{1}{2}\bigg{(}\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\tau_{\mathrm{sasv}})+\mathcal{L}_{\mathrm{BCE}}\bigg{)}.$ (7)
The algorithm then conducts a grid search (Line 12 in Algorithm 1) for a value of $\tau$ that gives the optimal threshold according to the a-DCF loss on the training data. The threshold value serves as the independent variable in Eq. (5) of the a-DCF loss. Using the a-DCF loss (or its combination with BCE loss) as the objective function during model training involves minimizing its value, which depends on the threshold value. This relationship necessitates finding the optimal threshold value that minimizes the a-DCF loss at each epoch which is achieved through grid search. This algorithm tracks the best model parameters and threshold as scored on development data, thus updating the lowest a-DCF found so far and its corresponding parameters. At the end, it returns the best threshold $\hat{\tau}_{*}$ and model parameters $\Theta_{*}$.
III Experimental Setup, Results and Discussion
We perform experiments on the ASVspoof2019 logical access (LA)  [3] dataset. The description of the dataset and experimental details are discussed in the following subsections.
III-A Dataset
The ASVspoof2019 LA subset consists of three parts, training, development, and evaluation. Each partition has a disjoint set of speakers. The average duration of the utterances is $2$ to $3$ seconds. The spoofed utterances are generated using $19$ VC and TTS algorithms ($6$ types of attacks in the training and the development sets, $13$ in the evaluation set).
III-B Experimental Setup
To demonstrate our proposed optimization scheme, we consider the DNN-based embedding fusion strategy ’Baseline2’ described in [7], that was used as a baseline in the SASV2022 challenge. Thereafter, many other studies follow similar approach to perform SASV [16, 17]. We aim to show whether our proposed technique brings any performance improvement or not. This model concatenates three embeddings ($\textbf{e}=\left[\textbf{e}_{\text{enr}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{cm}}\right]$): two speaker (ASV) embeddings from the enrolment ($\textbf{e}_{\text{enr}}^{\text{asv}}$) and test ($\textbf{e}_{\text{tst}}^{\text{asv}}$) utterances extracted using ECAPA-TDNN [18], and one spoofing (CM) embedding ($\textbf{e}_{\text{tst}}^{\text{cm}}$) extracted from the test utterance using AASIST [19]. The concatenated embeddings are fed to a DNN model that consists of three fully connected hidden layers with leaky ReLU activation functions. The layers contain $256$, $128$, and $64$ neurons, respectively. Unlike baseline, which contains two neurons with softmax, our model’s final layer contains a single neuron with a sigmoid activation function. This is because of our a-DCF optimization requires only one output score.
We consider four back-end models that are otherwise identical but optimized differently. These methods are:
  • S1: corresponds to the Baseline2 system provided by SASV challenge organizers (optimized with CE);
  • S2: optimized using soft a-DCF loss with, $\tau_{\mathrm{sasv}}=0.5$;
  • S3: optimized using soft a-DCF loss + BCE, with $\tau_{\mathrm{sasv}}=0.5$;
  • S4: optimized using soft a-DCF loss + BCE, including threshold optimization, as detailed in Algorithm 1.
The cost values $C_{\mathrm{miss}}^{\mathrm{tar}}$, $C_{\mathrm{fa}}^{\mathrm{non}}$, and $C_{\mathrm{fa}}^{\mathrm{spf}}$ in the soft a-DCF calculation are set as $1$, $10$, and $20$, respectively, while the prior values $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$ are set as $0.9$, $0.05$, and $0.05$. The higher relative cost value for spoof false alarms ($C_{\mathrm{fa}}^{\mathrm{spf}}=20$) is chosen based on the importance of spoofing robustness in security-sensitive domains, such as banking, where preventing unauthorized access is crucial. This reflects the emphasis to penalize false acceptance of spoof attempts to ensure the integrity and security of such systems. Additionally, the high target prior ($\pi_{\mathrm{tar}}=0.9$) indicates that the system is designed under the assumption that most interactions are legitimate. This is practical for applications where legitimate access attempts are frequent, and it minimizes the disruption to bonafide users by reducing the likelihood of false rejections. Except training, we considered the a-DCF metric given in Eq. 2 in all the experiments for performance evaluation.
TABLE I: Baseline system min a-DCF values in the development set
Batch size 24 48 96 192 384 768 1024
min a-DCF 0.1437 0.1308 0.1268 0.1346 0.1297 0.1251 0.1234
III-C Baseline (S1)
We firstly trained a baseline (S1) with cross entropy loss for various batch sizes, resulting to minimum a-DCF values shown in Table I. The results indicate that the best performance was achieved with a batch size of $1024$. We fix this value for the remainder of the experiments.
TABLE II: Comparison of S1 and S2 in development set. Batch size is selected as 1024.
System min a-DCF
S2 0.1355
S3 0.1182
III-D Proposed a-DCF metric optimization (S2, S3 and S4)
We next consider models S2 and S3, beginning with the hard-coded detection threshold of $\tau_{\mathrm{sasv}}=0.5$ in Eq. (5). Model selection is based on the lowest a-DCF value at a threshold of $0.5$ in the development set. The soft a-DCF only system (S2) attains minimum a-DCF of $0.1355$ and the combination of the soft a-DCF and BCE (S3) achieves minimum a-DCF of $0.1182$ as summarized in Table II.
TABLE III: Comparison of S1, S2, S3, and S4 in development and evaluation set. The results are given using 1024 as batch size.
Baseline (S1) S2 S3 S4
Dev 0.1234 0.1355 0.1182 0.1109
Eval 0.1445 0.2352 0.1398 0.1254
The results of S2 and S3 were based on hard-coded arbitrary detection threshold of $\tau_{\mathrm{sasv}}=0.5$. We now also optimize $\tau_{\mathrm{sasv}}$ using the complete procedure (Algorithm 1). Joint optimization of the DNN parameters and the threshold leads to noticeable improvements, as expected. Specifically, S4 achieved a minimum a-DCF of $0.1109$ on the development set and $0.1254$ on the evaluation set, outperforming S1, S2, and S3 as Table III indicates.
Figure 2: The plot illustrates the a-DCF values with respect to the threshold for the four different systems. The curves represent the a-DCF values for Baseline (S1), S2, S3, and S4 systems across various threshold values. Circular markers indicate the a-DCF values at a fixed threshold of $0.5$ for each system. Cross markers denote the minimum a-DCF values and their corresponding threshold points. The star marker represents the optimized threshold and the corresponding a-DCF value proposed for the S4 system.
Fig. 2 illustrates the a-DCF curves obtained for the four different systems (S1, S2, S3, and S4) as a function of the threshold. For systems S2 and S3, the a-DCF values at both the fixed threshold value of $0.5$ and their respective minimum points are depicted on the graph. For the S4 system, that includes threshold optimization, the a-DCF value corresponding to the optimized threshold is marked with a star. As can be seen from the graph (and supported by the results), the S4 model provides the lowest a-DCF value.
III-E Impact of a-DCF Parameter
Figure 3: Target vs Nontarget and Target vs Spoof DET curves for three different a-DCF parameter settings. The SV-EER values for Setting1 (represented by the solid blue line), Setting1 (soft a-DCF + BCE) (represented by the dotted blue line), Setting2, and Setting3 are 13.97%, 8.44%, 7.52%, and 49.13%, respectively.
Our final analysis addresses the impact of the a-DCF parameters. To this end, we introduced three different set of a-DCF parameters as:
  • Setting1 (soft a-DCF) contains only a-DCF loss. In this case, the cost values $C_{\mathrm{miss}}^{\mathrm{tar}}$, $C_{\mathrm{fa}}^{\mathrm{non}}$, and $C_{\mathrm{fa}}^{\mathrm{spf}}$ are set to $1$, $10$, and $20$, respectively. The prior values $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$ are set to $0.9$, $0.05$, and $0.05$.
  • Setting1 (soft a-DCF + BCE) contains both a-DCF and BCE losses. The cost and the prior values are the same as in the previous case.
  • Setting2 (soft a-DCF) contains only a-DCF loss. In this case, the cost values $C_{\mathrm{miss}}^{\mathrm{tar}}$, $C_{\mathrm{fa}}^{\mathrm{non}}$, and $C_{\mathrm{fa}}^{\mathrm{spf}}$ are all set to $1$. The prior values $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$ are set to $0.5$, $0.5$, and 00, respectively.
  • Setting3 (soft a-DCF) contains only a-DCF loss. The same cost values are used as in the previous setting, while $0.5$, 00 and $0.5$ priors are used for $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$, respectively.
The threshold optimization method was included for all these cases. Fig. 3 shows target vs. nontarget and target vs. spoof DET profiles for three different sets of a-DCF parameters. The figure clearly indicates that the choice of a-DCF parameters significantly impacts both DET profiles. Comparing setting1’s soft a-DCF alone with the soft a-DCF combined with BCE, the former generally has a lower miss rate. For setting2 and setting3, changes in the priors cause notable differences in the DET curves. For example, setting $\pi_{\mathrm{spf}}$ to 00 (setting2), assuming no spoof non-targets, shows the best performance for target vs. nontarget trials across all operating points. However, when the non-target prior is set to 00 (setting3), the performance for target vs. nontarget trials is found to be inferior, as expected. These trends are reversed for target vs. spoof trials.
IV Conclusion
In this study, we demonstrated that a spoofing-aware speaker verification system can be effectively optimized for the new a-DCF metric. By incorporating both the a-DCF and binary cross-entropy losses in the optimization process, we fine-tuned the network weights, which led to improved performance over the baseline system. Importantly, the usage of a novel threshold optimization technique further enhanced the system’s capabilities, resulting in a significant improvement in a-DCF values. These findings highlight the potential for adjusting ASV systems to achieve an optimal balance between user convenience and security against to evolving spoofing attacks.

### v2

@@
 Abstract
-Automatic speaker verification (ASV) systems are vulnerable to spoofing attacks such as text-to-speech.
-In this study, we propose a novel spoofing-robust ASV back-end classifier, optimized directly for the recently introduced, architecture-agnostic detection cost function (a-DCF).
-We combine a-DCF and binary cross-entropy (BCE) losses to optimize the network weights, combined by a novel, straightforward detection threshold optimization technique.
-Experiments on the ASVspoof2019 database demonstrate considerable improvement over the baseline optimized using BCE only (from minimum a-DCF of $0.1445$ to $0.1254$), representing $13\%$ relative improvement.
-These initial promising results demonstrate that it is possible to adjust an ASV system to find appropriate balance across the contradicting aims of user convenience and security against adversaries.
+Automatic speaker verification (ASV) systems are vulnerable to spoofing attacks.
+We propose a spoofing-robust ASV system optimized directly for the recently introduced architecture-agnostic detection cost function (a-DCF), which allows targeting a desired trade-off between the contradicting aims of user convenience and robustness to spoofing.
+We combine a-DCF and binary cross-entropy (BCE) with a novel straightforward threshold optimization technique.
+Our results with an embedding fusion system on ASVspoof2019 data demonstrate relative improvement of $13\%$ over a system trained using BCE only (from minimum a-DCF of $0.1445$ to $0.1254$).
+Using an alternative non-linear score fusion approach provides relative improvement of $43\%$ (from minimum a-DCF of $0.0508$ to $0.0289$).
 Index Terms:
@@
 Automatic speaker verification (ASV) seeks to verify a claimed identity based on speech evidence [1].
-Despite being a matured biometric technology, ASV is susceptible to spoofing through replay [2], text-to-speech [3] and other adversaries.
+Despite being a matured biometric technology, ASV is susceptible to spoofing through replay [2], text-to-speech [3] and other attacks.
 To counter these risks, various countermeasures (CMs) or presentation attack detectors (PADs) [4], have been proposed for detecting spoofing.
-Typically CMs are developed in isolation from ASV, and the two subsystems are combined to enhance robustness of the biometric system against adversaries.
+Typically CMs are developed in isolation from ASV, and the two subsystems are combined to enhance security.
 The common strategies include tandem (cascade) combination [5] and fusion of detection scores [6] or embeddings [7, 8].
-Besides systems constructed from combination of speaker and spoof detection subsystems, other studies consider deep end-to-end networks optimized to predict the decision directly from the speech inputs [9, 10].
-Whether realized by combining two subsystems (ASV and CM), or by training a monolithic end-to-end model, a key consideration is how to evaluate and optimize the complete biometric authentication system.
-Any such system must trade-off between misses (false rejections) and false alarms (false acceptances)—proxies for user (in)convenience and security, respectively.
-The so-called tandem detection cost function (t-DCF) [5] can be used to evaluate the performance of cascaded combinations of ASV and CM under a Bayes-risk framework.
-Nonetheless, as t-DCF requires two sets of detection scores (and two detection thresholds) corresponding to the ASV and the CM subsystems, it is inapplicable for performance assessment of other types of architectures.
-To address these limitations, [11] proposed recently an architecture-agnostic detection cost function (a-DCF) applicable to a wider variety of system architectures.
-Unlike t-DCF, a-DCF requires a single detection score and a single threshold value only.
-This new metric is the primary metric in one of the two main tracks in the currently running ASVspoof 5 challenge [^1] https://www.asvspoof.org/ (referred June 28, 2024).
-For the first time, we propose a spoofing-robust ASV back-end that is directly optimized for the new a-DCF metric (as shown in Fig.
+Besides systems constructed from combination of speaker and spoof detection subsystems, monolithic end-to-end solutions [9, 10], [11] have also been considered.
+Whether realized as a combination of the ASV and CM subsystems, or an end-to-end model, a key consideration is how to evaluate and optimize the complete system, taking into account the trade-off between misses (false rejections) and false alarms (false acceptances)—proxies for user (in)convenience and security, respectively.
+The so-called tandem detection cost function (t-DCF) [5] can be used to evaluate CM-ASV cascades under a Bayes-risk framework.
+Nonetheless, t-DCF requires two sets of detection scores (and two thresholds) corresponding to the two subsystems, making it inapplicable for evaluation of any other type of architecture.
+To address this limitation, [12] proposed recently an architecture-agnostic detection cost function (a-DCF) applicable to a wider variety of architectures.
+Unlike t-DCF, a-DCF requires only one set of detection scores and one detection threshold.
+This new metric was the primary metric in one of the two main tracks in the ASVspoof 5 challenge [13].
+Figure 1: The proposed a-DCF optimization for spoofing-robust speaker verification uses embeddings or scores from enrollment ($\mathcal{O}_{\text{enr}}$) and test ($\mathcal{O}_{\text{tst}}$) inputs to give scores $g(x)$ for target, non-target, and spoof classes.
+Error rates $\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}$, $\hat{P}_{\mathrm{fa}}^{\mathrm{non}}$, and $\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}$, computed using threshold $\tau_{\text{sasv}}$, form the soft a-DCF loss for backpropagation to optimize class-specific detection accuracy.
+For the first time, we address spoofing-robust ASV directly optimized for the a-DCF metric (Fig.
 1).
-As an evaluation metric based on hard error counting, the original a-DCF metric is non-differentiable and cannot be directly optimized using gradient descent schemes.
+As a metric based on hard error counting, a-DCF is non-differentiable and hence cannot be directly optimized using gradient descent schemes.
 Luckily, hard error counts can be “softened” (made differentiable).
@@
 By integrating a differentiable version of the a-DCF into model training, we aim to enhance the performance and robustness of speaker verification systems against spoofing attacks, ultimately contributing to the advancement of secure and reliable biometric authentication technologies.
-Besides integrating the a-DCF metric into model training, the threshold value is also jointly optimized during model training in the proposed method.
-Figure 1: Proposed a-DCF optimization method for spoofing-robust speaker verification.
-The model $g(x)$ takes embeddings as input and outputs score distributions, from which error rates $\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}$, $\hat{P}_{\mathrm{fa}}^{\mathrm{non}}$, and $\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}$ are calculated.
-These error rates are then used to compute soft a-DCF, which serve as the loss function for model optimization through backpropagation, ensuring improved performance by minimizing the a-DCF.
+Besides integrating the a-DCF metric into model training, the threshold ($\tau_{\mathrm{sasv}}$) is jointly optimized with the model parameters in the proposed method.
 II Detection Cost Functions
-Conventional detection cost functions [13] are evaluation metrics used to assess the performance of binary classification systems, such as speaker verification and anti-spoofing.
-They focus on balancing the trade-off between false alarms and misses, assigning specific costs to each type of error.
-A binary classifier can be defined as a function $g:\mathcal{X}\to\rm I\!R$ that assigns a detection score $g(x)$ for a given input $x\in\mathcal{X}$.
-This score then can be converted into predicted binary class label $\hat{y}\in\{0,1\}$ by comparing it with a pre-set threshold $\tau$ by $\hat{y}=\mathrm{I}(g(x)>\tau)$.
-Here, $\mathrm{I}$ is the indicator function which returns $1$ if the condition is true (null hypothesis accepted) and 00 otherwise.
-In conventional ASV, $x$ consists of a pair of enrollment and test utterances (presented typically as a pair of speaker embeddings), and the null hypothesis is that the same speaker is present in both.
-In our spoofing-aware ASV models, $x$ additionally contains spoof embedding extracted from the test utterance by a CM model.
+Detection cost functions [15] are evaluation metrics used to assess the performance of binary classifiers, such as speaker verification and anti-spoofing systems.
+A binary classifier can be defined as a function $g:\mathcal{X}\to\mathbb{R}$ that assigns a detection score $g(x)$ for an input $x\in\mathcal{X}$, converted subsequently into predicted label $y\in\{0,1\}$ by comparing it with a threshold $\tau$, i.e.
+$y=\mathrm{I}(g(x)>\tau)$.
+Here, $\mathrm{I}(\cdot)$ is the indicator function which returns $1$ if the condition is true (null hypothesis accepted) and 00 otherwise.
+In ASV, $x=(\mathcal{O}_{\text{enr}},\mathcal{O}_{\text{tst}})$ consists of a pair of enrollment $(\mathcal{O}_{\text{enr}})$ and test $(\mathcal{O}_{\text{tst}})$ utterances.
+The null hypothesis is that the same speaker produced both utterances.
+To develop a spoofing-aware speaker verification (SASV) system, we represent $x$ either as a triplet of multivariate embeddings $(\boldsymbol{\mathrm{e}}_{\text{enr}}^{\text{asv}},\boldsymbol{\mathrm{e}}_{\text{tst}}^{\text{asv}},\boldsymbol{\mathrm{e}}_{\text{tst}}^{\text{cm}})$, or as a pair of scalars $(s_{\text{asv}},s_{\text{cm}})$.
+In the former case, the two speaker embeddings $(\boldsymbol{\mathrm{e}}_{\text{enr}}^{\text{asv}},\boldsymbol{\mathrm{e}}_{\text{tst}}^{\text{asv}})$ extracted using an ASV model are augmented by a spoof embedding $\boldsymbol{\mathrm{e}}_{\text{tst}}^{\text{cm}}$ extracted using a CM model.
+The latter case, consists of a pair of ASV and CM scores We demonstrate our back-end optimization for both types of input representations.
 II-A DCF and a-DCF
-Detection Cost Function (DCF) [13] is used for performance assessment of conventional ASV systems. It is defined as:
-$\mathrm{DCF(\tau_{asv})}=C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{asv}})+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{asv}}),$ (1)
-where $C_{\mathrm{miss}}^{\mathrm{tar}}$ is the cost of a miss, $C_{\mathrm{fa}}^{\mathrm{non}}$ is the cost of a false alarm, $P_{\mathrm{miss}}^{\mathrm{tar}}$ is the miss rate, $P_{\mathrm{fa}}^{\mathrm{non}}$ is false alarm rate, $\pi_{\mathrm{tar}}$ is the prior probability of the target and $\pi_{\mathrm{non}}=1-\pi_{\mathrm{tar}}$ is the prior probability of nontarget, and $\tau_{\mathrm{asv}}$ is the detection threshold.
+Detection Cost Function (DCF) [15] is used for performance assessment of conventional ASV systems.
+It considers consequences of both false alarms (FAs) and misses through a cost model.
+The DCF is formally defined as,
+$\footnotesize\mathrm{DCF(\tau_{asv})}=C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{asv}})+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{asv}}),$ (1)
+where $C_{\mathrm{miss}}^{\mathrm{tar}}$ and $C_{\mathrm{fa}}^{\mathrm{non}}$ are the costs of a miss and FA, respectively, $P_{\mathrm{miss}}^{\mathrm{tar}}$ and, $P_{\mathrm{fa}}^{\mathrm{non}}$ are the miss and FA rates, $\pi_{\mathrm{tar}}$ and $\pi_{\mathrm{non}}=1-\pi_{\mathrm{tar}}$ represent target and nontarget prior probabilities, and $\tau_{\mathrm{asv}}$ is the detection threshold.
 The DCF in Eq.
-(1) allows one to set the cost parameters according to the user’s application specification.
-The architecture-agnostic detection cost function (a-DCF) [11] is a performance metric introduced for spoofing-aware speaker verification (SASV) task.
-It extends upon the DCF in Eq.
-(1) by including a third class (spoofing attack):
+(1) allows tuning of cost parameters to specific application needs.
+Architecture-agnostic detection cost function (a-DCF) [12] is used for assessing spoofing-robust ASV systems. It extends the DCF by including a third class (spoofing attack):
 $\mathrm{a\!-\!DCF}(\tau_{\mathrm{sasv}})=$ $\;C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ (2)
@@
 $+C_{\mathrm{fa}}^{\mathrm{spf}}\cdot\pi_{\mathrm{spf}}\cdot P_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}}),$
-where $P_{\mathrm{fa}}^{\mathrm{spf}}$ is the false alarm for spoof trials, $\pi_{\mathrm{spf}}$ is the prior probability of spoofing, $C_{\mathrm{fa}}^{\mathrm{spf}}$ is the respective costs for spoof false alarms.
-$\tau_{\mathrm{sasv}}$ is the detection threshold for spoofing-aware speaker verification (SASV) system.
-The miss rate ($P_{\mathrm{miss}}^{\mathrm{tar}}$) and the false alarm rates ($P_{\mathrm{fa}}^{\mathrm{non}}$, $P_{\mathrm{fa}}^{\mathrm{spf}}$) in the a-DCF metric are computed as follows:
+where $P_{\mathrm{fa}}^{\mathrm{spf}}$, $\pi_{\mathrm{spf}}$, $C_{\mathrm{fa}}^{\mathrm{spf}}$ and $\tau_{\mathrm{sasv}}$ are the FA rate of spoofs, prior of spoofing, cost of spoof FA, and the threshold, respectively.
+The three error rates in (2) are computed as follows:
 $P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{tar}}}\sum_{x\in\mathrm{tar}}\mathrm{I}(g(x)\leq\tau_{\mathrm{sasv}})$ (3)
@@
 $P_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\text{sasv}})$ $=\frac{1}{N_{\mathrm{spf}}}\sum_{x\in\mathrm{spf}}\mathrm{I}(g(x)>\tau_{\mathrm{sasv}}),$
-where $N_{\mathrm{tar}}$, $N_{\mathrm{non}}$, and $N_{\mathrm{spf}}$ represent the number of target, non-target, and spoof class trials, respectively.
+where $\mathrm{tar}$, $\mathrm{non}$ and $\mathrm{spf}$ denote the sets of target, non-target, and spoof trials, respectively, with their counts denoted by $N_{\bullet}$.
 II-B Differentiable a-DCF Metric
-The indicator function used to compute false alarm and miss rates in Eq.
+The indicator function in Eq.
 (3) is non-differentiable.
-Therefore, the a-DCF metric can not be directly used as loss function to optimize an SASV system.
-In this study, we propose a differentiable (soft) version of a-DCF suitable for direct minimization using gradient descent.
-Similar approaches have been successfully applied to optimize DCF for text-dependent speaker verification [14] and t-DCF for the constrained class of tandem systems [15], but not the a-DCF.
-The a-DCF metric provides a more holistic evaluation by integrating the costs associated with speaker verification.
-This approach simplifies the benchmarking process for different ASV system architectures, as it eliminates the need for separate metrics for each component.
-By focusing on a single metric, the a-DCF metric promotes development of integrated systems that are optimized for spoofing-aware speaker verification.
+Therefore, the a-DCF metric can not be directly used as loss function in gradient search methods.
+In this work, we propose a differentiable (soft) version of a-DCF suitable for this purpose.
+Similar approaches have been successfully used to optimize DCF for text-dependent speaker verification [16] and t-DCF for tandem architectures [14].
+Compared with t-DCF, applicable only to the unnecessarily restrictive class of tandem systems, a-DCF is applicable to arbitrary SASV systems.
+Refer to [12] for further elaboration.
 The non-differentiable components in a-DCF metric are the miss rate ($P_{\text{miss}}$) and the two false alarm rates ($P_{\text{fa}}^{\text{non}}$, $P_{\text{fa}}^{\text{spf}}$).
@@
 $\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{spf}}}\sum_{x\in\mathrm{spf}}\sigma\left(g(x)-\tau_{\mathrm{sasv}}\right),$
-where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function. Essentially, the error rates are approximated based on their distance from the threshold. The soft a-DCF loss is now defined as:
-$\mathcal{L}_{\mathrm{a\!-\!DCF}}^{\mathrm{soft}}(\tau_{\mathrm{sasv}})=$ $\;C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ (5)
-$+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot\hat{P}_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$
-$+C_{\mathrm{fa}}^{\mathrm{spf}}\cdot\pi_{\mathrm{spf}}\cdot\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}}).$
+where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function.
+Essentially, the error rates are approximated based on their distance from the threshold.
+The soft a-DCF loss is obtained by replacing the error rate terms in Eq.
+(3) with the expressions from Eq.
+(4), applied within the original a-DCF formula in Eq.
+(2).
 II-C Optimization of soft a-DCF
@@
 1:Input: Train data $\mathcal{D}_{\mathrm{trn}}$, Development data $\mathcal{D}_{\mathrm{dev}}$, Batch Size $B$, Number of Epochs $N$
-2:Output: Best threshold $\hat{\tau}_{*}$, Model Parameters $\Theta_{*}$
+2:Output: Best threshold $\tau_{\mathrm{sasv}}$, Model Parameters $\Theta_{*}$
 3:Initialize: $\mathrm{min\_a\!-\!DCF}\leftarrow\infty$, $\tau\leftarrow 0.5$
-4:for epoch $=1$ to $N$ do
+4:for i $=1$ to $N$ do
 5:    // optimize DNN weights
@@
 15:         $\mathrm{min\_a\!-\!DCF}\leftarrow\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\tau_{i+1})$
-16:         $\hat{\tau}\leftarrow\tau_{i+1}$
-17:    end if
-18:end for
-19:return $\hat{\tau}_{*}$, $\Theta_{*}$
-The algorithm 1 presents the training loop of a neural network towards its optimization, with both soft a-DCF ($\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(X;\Theta_{i},\tau_{i})$ ) and BCE loss ( $\mathcal{L}_{\mathrm{BCE}}(X;\Theta_{i})$ ) functions.
-The BCE loss is the binary cross entropy loss defined as:
-$\mathcal{L}_{\mathrm{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}[y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})]$ (6)
-where $N$ represents the number of training samples.
-Essentially, the soft a-DCF loss is a loss function of application dependent a-DCF metric and useful for fine-tuning the system towards a desired miss-false alarm balance.
-The BCE loss, in turn, helps to separate target from the pool of the two negative classes.
-For these reasons, we use the combination of these two losses.
+16:         $\tau_{\mathrm{sasv}}\leftarrow\tau_{i+1}$
+17:         $\Theta_{*}\leftarrow\Theta$
+18:    end if
+19:end for
+20:return $\tau_{\mathrm{sasv}}$, $\Theta_{*}$
+The algorithm 1 presents the optimization loop of a generic model through a combination of soft a-DCF ($\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(X;\Theta_{i},\tau_{i})$) and BCE ($\mathcal{L}_{\mathrm{BCE}}(X;\Theta_{i})$) losses.
+The former is defined as,
+$\footnotesize\mathcal{L}_{\mathrm{BCE}}=-\frac{1}{N}\sum_{j=1}^{N}[y_{j}\log(g(x_{j}))+(1-y_{j})\log(1-g(x_{j}))],$ (5)
+where $N$ is the number of training samples and $g(x_{j})$ is the model output (score).
 The algorithm initializes a minimum a-DCF to be infinity and a fixed threshold $\tau$.
 Then, for each epoch, it optimizes the model weights by calculating the combined loss $\mathcal{J}(\Theta_{i},\tau_{i})$ from soft a-DCF and BCE losses and updating the weights of the neural network.
-This combined loss is defined as:
-$J(\Theta_{i},\tau_{i})=\frac{1}{2}\bigg{(}\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\tau_{\mathrm{sasv}})+\mathcal{L}_{\mathrm{BCE}}\bigg{)}.$ (7)
+This combined loss is defined as Line 7 in Algorithm 1.
 The algorithm then conducts a grid search (Line 12 in Algorithm 1) for a value of $\tau$ that gives the optimal threshold according to the a-DCF loss on the training data.
@@
 This algorithm tracks the best model parameters and threshold as scored on development data, thus updating the lowest a-DCF found so far and its corresponding parameters.
-At the end, it returns the best threshold $\hat{\tau}_{*}$ and model parameters $\Theta_{*}$.
+At the end, it returns the best threshold $\tau_{\mathrm{sasv}}$
+To elaborate on the combined loss in Line 7, assume that the empirical proportion of the training data from each class (target, nontarget, spoof) are equal ($\sim\frac{1}{3}$) in each minibatch.
+Expanding the combined loss $\mathcal{J}(\Theta_{i},\tau_{i})$, after some algebraic manipulation, can be shown to give
+$\mathcal{J}\propto$ $\frac{1}{N_{\text{tar}}}\sum_{x\in\text{tar}}\left[A\sigma\left(\tau_{\text{sasv}}-g(x)\right)-\log g(x)\right]$ (6)
+$+\frac{1}{N_{\text{non}}}\sum_{x\in\text{non}}\left[B\sigma\left(g(x)-\tau_{\text{sasv}}\right)-\log(1-g(x))\right]$
+$+\frac{1}{N_{\text{spf}}}\sum_{x\in\text{spf}}\left[C\sigma\left(g(x)-\tau_{\text{sasv}}\right)-\log(1-g(x))\right],$
+where $A$, $B$ and $C$ are constants implied by the a-DCF parameters111 In specific, $A:=3C_{\text{tar}}^{\text{miss}}\pi_{\text{tar}}$, $B:=3C_{\text{non}}^{\text{fa}}\pi_{\text{non}}$, and $C:=3C_{\text{spf}}^{\text{fa}}\pi_{\text{spf}}$.
+and where $\propto$ signifies omission of unimportant scaling constants.
+The combined loss is balanced average of the three class-conditional loss terms, with the per-instance losses enclosed within the brackets.
+The relative contribution of the two types of terms depends on the a-DCF parameters.
+While the sigmoid (from a-DCF) and the log (from BCE) terms are tied through the score $g(x)$, $\tau_{\text{sasv}}$ is present only in the a-DCF terms—but tied across the class-conditional losses.
+As we will demonstrate below, the interplay of the two losses is important in optimization.
+While BCE focuses on class separation (discrimination), a-DCF takes into account the costs of actual decisions by including the decision threshold.
+Together, BCE and a-DCF form a complementary loss function that not only establishes a robust classification boundary but also adapts the model to prioritize specific classes based on their operational relevance and associated costs.
 III Experimental Setup, Results and Discussion
-We perform experiments on the ASVspoof2019 logical access (LA)  [3] dataset. The description of the dataset and experimental details are discussed in the following subsections.
 III-A Dataset
-The ASVspoof2019 LA subset consists of three parts, training, development, and evaluation.
-Each partition has a disjoint set of speakers.
-The average duration of the utterances is $2$ to $3$ seconds.
-The spoofed utterances are generated using $19$ VC and TTS algorithms ($6$ types of attacks in the training and the development sets, $13$ in the evaluation set).
+The SASV experiments are conducted on the ASVspoof2019 logical access (LA) dataset [3], which consists of training, testing, and evaluation subsets, with disjoint speaker identity.
+This dataset includes both bonafide and spoofed utterances, with spoofed signals generated using various voice conversion (VC) and text-to-speech (TTS) algorithms.
+A detailed dataset description is available at [3].
 III-B Experimental Setup
-To demonstrate our proposed optimization scheme, we consider the DNN-based embedding fusion strategy ’Baseline2’ described in [7], that was used as a baseline in the SASV2022 challenge.
+To demonstrate our proposed optimization scheme, we consider the DNN-based embedding fusion strategy ’Baseline2’ described in [7], used as a baseline in the SASV2022 challenge.
 Thereafter, many other studies follow similar approach to perform SASV [17, 18].
-We aim to show whether our proposed technique brings any performance improvement or not.
-This model concatenates three embeddings ($\textbf{e}=\left[\textbf{e}_{\text{enr}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{cm}}\right]$): two speaker (ASV) embeddings from the enrolment ($\textbf{e}_{\text{enr}}^{\text{asv}}$) and test ($\textbf{e}_{\text{tst}}^{\text{asv}}$) utterances extracted using ECAPA-TDNN [18], and one spoofing (CM) embedding ($\textbf{e}_{\text{tst}}^{\text{cm}}$) extracted from the test utterance using AASIST [19].
-The concatenated embeddings are fed to a DNN model that consists of three fully connected hidden layers with leaky ReLU activation functions.
-The layers contain $256$, $128$, and $64$ neurons, respectively.
-Unlike baseline, which contains two neurons with softmax, our model’s final layer contains a single neuron with a sigmoid activation function.
-This is because of our a-DCF optimization requires only one output score.
-We consider four back-end models that are otherwise identical but optimized differently. These methods are:
+The two speaker embeddings $(\textbf{e}_{\text{enr}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{asv}})$ are extracted through ECAPA-TDNN [19] and the spoof embedding ($\textbf{e}_{\text{tst}}^{\text{cm}}$) through AASIST [20].
+The three concatenated embeddings are fed to a DNN model that consists of three fully connected hidden layers of $256$, $128$, and $64$ neurons with leaky ReLU activation functions.
+Unlike the baseline’s two-neuron softmax output, our model’s final layer has a single neuron with a sigmoid activation, as a-DCF optimization requires only one output score.
+We consider four back-end models with identical architectures, optimized differently. They are:
   • S1: corresponds to the Baseline2 system provided by SASV challenge organizers (optimized with CE);
-  • S2: optimized using soft a-DCF loss with, $\tau_{\mathrm{sasv}}=0.5$;
-  • S3: optimized using soft a-DCF loss + BCE, with $\tau_{\mathrm{sasv}}=0.5$;
-  • S4: optimized using soft a-DCF loss + BCE, including threshold optimization, as detailed in Algorithm 1.
+  • S2: optimized using soft a-DCF, with $\tau_{\mathrm{sasv}}=0.5$;
+  • S3: optimized using soft a-DCF + BCE, with $\tau_{\mathrm{sasv}}=0.5$;
+  • S4: optimized using soft a-DCF + BCE, including threshold optimization, as detailed in Algorithm 1.
+In addition to these embedding fusion methods, we proposed a-DCF optimization of a trainable nonlinear score fusion method [21] (detailed below).
 The cost values $C_{\mathrm{miss}}^{\mathrm{tar}}$, $C_{\mathrm{fa}}^{\mathrm{non}}$, and $C_{\mathrm{fa}}^{\mathrm{spf}}$ in the soft a-DCF calculation are set as $1$, $10$, and $20$, respectively, while the prior values $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$ are set as $0.9$, $0.05$, and $0.05$.
@@
 This reflects the emphasis to penalize false acceptance of spoof attempts to ensure the integrity and security of such systems.
-Additionally, the high target prior ($\pi_{\mathrm{tar}}=0.9$) indicates that the system is designed under the assumption that most interactions are legitimate.
-This is practical for applications where legitimate access attempts are frequent, and it minimizes the disruption to bonafide users by reducing the likelihood of false rejections.
-Except training, we considered the a-DCF metric given in Eq.
-2 in all the experiments for performance evaluation.
+Additionally, the high target prior ($\pi_{\mathrm{tar}}=0.9$) aligns with an assumption that most interactions are legitimate.
+The a-DCF metric in (2) is used for performance evaluation.
+Additionally, equal error rate (EER) is reported in selected cases.
+We use SV-EER (speaker verification EER) to assess the system’s ability to differentiate between target and non-target samples, and SPF-EER (spoofing EER) to assess ability to differentiate between bonafide targets and spoofing attacks.
 TABLE I: Baseline system min a-DCF values in the development set
-Batch size 24 48 96 192 384 768 1024
-min a-DCF 0.1437 0.1308 0.1268 0.1346 0.1297 0.1251 0.1234
+Batch size 24 48 96 192 384 768 1024 2048
+min a-DCF 0.1437 0.1308 0.1268 0.1346 0.1297 0.1251 0.1234 0.1243
 III-C Baseline (S1)
@@
 We fix this value for the remainder of the experiments.
-TABLE II: Comparison of S1 and S2 in development set. Batch size is selected as 1024.
-System min a-DCF
-S2 0.1355
-S3 0.1182
 III-D Proposed a-DCF metric optimization (S2, S3 and S4)
-We next consider models S2 and S3, beginning with the hard-coded detection threshold of $\tau_{\mathrm{sasv}}=0.5$ in Eq.
-(5).
+We next consider models S2 and S3, beginning with the hard-coded detection threshold of $\tau_{\mathrm{sasv}}=0.5$ in the soft a-DCF loss equation.
 Model selection is based on the lowest a-DCF value at a threshold of $0.5$ in the development set.
-The soft a-DCF only system (S2) attains minimum a-DCF of $0.1355$ and the combination of the soft a-DCF and BCE (S3) achieves minimum a-DCF of $0.1182$ as summarized in Table II.
-TABLE III: Comparison of S1, S2, S3, and S4 in development and evaluation set. The results are given using 1024 as batch size.
-Baseline (S1) S2 S3 S4
-Dev 0.1234 0.1355 0.1182 0.1109
-Eval 0.1445 0.2352 0.1398 0.1254
-The results of S2 and S3 were based on hard-coded arbitrary detection threshold of $\tau_{\mathrm{sasv}}=0.5$.
+The soft a-DCF only system (S2) attains a minimum a-DCF of $0.1355$, with an SV-EER of $9.76\%$ and an SPF-EER of $0.14$ in the development set, while in the evaluation set, it reaches a minimum a-DCF of $0.2352$, an SV-EER of $17.28\%$ and an SPF-EER of $1.02\%$.
+In contrast, the combination of the soft a-DCF and BCE (S3) achieves a minimum a-DCF of $0.1182$, with an SV-EER of $8.29\%$ and an SPF-EER of $0.10\%$ in the development set, and in the evaluation set, it attains a minimum a-DCF of $0.1398$, an SV-EER of $9.59\%$ and an SPF-EER of $0.63\%$, indicating improved accuracy.
+TABLE II: Comparison of S1, S2, S3, and S4 in development and evaluation set. The results are given using 1024 as batch size.
+System a-DCF SV-EER SPF-EER
+Dev Eval Dev Eval Dev Eval
+Baseline (S1) 0.1234 0.1445 8.36 9.86 0.07 0.63
+S2 0.1355 0.2352 9.76 17.28 0.14 1.02
+S3 0.1182 0.1398 8.29 9.59 0.10 0.63
+S4 0.1109 0.1254 7.75 8.44 0.08 0.61
+The results of S2 and S3 were based on the hard-coded arbitrary detection threshold of $\tau_{\mathrm{sasv}}=0.5$.
 We now also optimize $\tau_{\mathrm{sasv}}$ using the complete procedure (Algorithm 1).
 Joint optimization of the DNN parameters and the threshold leads to noticeable improvements, as expected.
-Specifically, S4 achieved a minimum a-DCF of $0.1109$ on the development set and $0.1254$ on the evaluation set, outperforming S1, S2, and S3 as Table III indicates.
-Figure 2: The plot illustrates the a-DCF values with respect to the threshold for the four different systems.
-The curves represent the a-DCF values for Baseline (S1), S2, S3, and S4 systems across various threshold values.
-Circular markers indicate the a-DCF values at a fixed threshold of $0.5$ for each system.
-Cross markers denote the minimum a-DCF values and their corresponding threshold points.
-The star marker represents the optimized threshold and the corresponding a-DCF value proposed for the S4 system.
+Specifically, S4 achieved a minimum a-DCF of $0.1109$, with an SV-EER of $7.75\%$ and an SPF-EER of $0.08$ on the development set, and in the evaluation set, it reached a minimum a-DCF of $0.1254$, an SV-EER of $8.44\%$ and an SPF-EER of $0.61\%$, outperforming S1, S2, and S3, as Table II indicates.
+Figure 2: a-DCF values:Baseline (S1), S2, S3, and S4 across different thresholds.
 Fig.
-2 illustrates the a-DCF curves obtained for the four different systems (S1, S2, S3, and S4) as a function of the threshold.
-For systems S2 and S3, the a-DCF values at both the fixed threshold value of $0.5$ and their respective minimum points are depicted on the graph.
-For the S4 system, that includes threshold optimization, the a-DCF value corresponding to the optimized threshold is marked with a star.
-As can be seen from the graph (and supported by the results), the S4 model provides the lowest a-DCF value.
+2 presents the a-DCF curves for four systems (S1, S2, S3, and S4) across varying thresholds.
+For S2 and S3, both the a-DCF values at the fixed threshold of $0.5$ and the minimum points are indicated.
+The S4 system, which incorporates threshold optimization, has its optimized a-DCF value marked with a star.
+As seen in the graph, S4 yields the lowest a-DCF value.
 III-E Impact of a-DCF Parameter
-Figure 3: Target vs Nontarget and Target vs Spoof DET curves for three different a-DCF parameter settings.
-The SV-EER values for Setting1 (represented by the solid blue line), Setting1 (soft a-DCF + BCE) (represented by the dotted blue line), Setting2, and Setting3 are 13.97%, 8.44%, 7.52%, and 49.13%, respectively.
-Our final analysis addresses the impact of the a-DCF parameters. To this end, we introduced three different set of a-DCF parameters as:
+Figure 3: Target vs.
+Nontarget and Target vs.
+Spoof DET curves for three a-DCF parameter settings.
+SV-EER values are 13.97% for Setting1 (solid blue), 8.44% for Setting1 (soft a-DCF + BCE) (dotted blue), 7.52% for Setting2, and 49.13% for Setting3.
+Our final analysis addresses the impact of the a-DCF parameters. To this end, we introduced three different sets of a-DCF parameters as:
   • Setting1 (soft a-DCF) contains only a-DCF loss.
@@
 spoof trials.
+III-F a-DCF Loss in Score Fusion
+In this section, we demonstrated SASV performance using score fusion (SF) optimized with soft a-DCF.
+As reported in [21], nonlinear calibrated score fusion (NL-Cal-SF) performs better than both linear score sum (L-SF) and linear calibrated score sum (L-Cal-SF) fusion.
+Building on this, we applied soft a-DCF (aDCF-NL-Cal-SF) optimization with nonlinear calibrated score fusion.
+Performance metrics and trainable parameter counts are provided in Table III.
+The results indicate that our proposed SF approach improves min a-DCF from $0.0508$ to $0.0289$ compared to nonlinear calibrated SF.
+Additionally, we see that SF achieves better SASV performance compared with embedding fusion, despite using significantly fewer parameters ($4$ compared to $180736$).
+TABLE III: Comparison of SASV systems with proposed score fusion.
+SASV System #Trainable Params min a-DCF
+Embedding Fusion 180736 0.1234
+L-SF 0 0.5311
+L-Cal-SF 4 0.0648
+NL-Cal-SF 4 0.0508
+aDCF-NL-Cal-SF 4 0.0289
 IV Conclusion
-In this study, we demonstrated that a spoofing-aware speaker verification system can be effectively optimized for the new a-DCF metric.
-By incorporating both the a-DCF and binary cross-entropy losses in the optimization process, we fine-tuned the network weights, which led to improved performance over the baseline system.
-Importantly, the usage of a novel threshold optimization technique further enhanced the system’s capabilities, resulting in a significant improvement in a-DCF values.
+In this study, we demonstrated that an SASV system either embedding or score fusion can be effectively optimized for the new a-DCF metric.
+Incorporating both a-DCF and BCE losses in the optimization improved performance for embedding fusion over the baseline system.
+Importantly, the usage of a novel threshold optimization technique further enhanced the system’s capabilities, resulting in improvement in a-DCF values.
 These findings highlight the potential for adjusting ASV systems to achieve an optimal balance between user convenience and security against to evolving spoofing attacks.